Question 1- Find 4 consecutive integers whose sum is twice the cube of 5.

Answer by CubeyThePenguin(3113)

x + (x+1) + (x+2) + (x+3) = 2(5^3)
4x + 6 = 250
4x = 244
x = 61

integers: 61, 62, 63, 64